# Hard rigid rods on Husimi trees

### v1

@@ Appendix A Solution of the model on the Husimi lattice built with triangles
We may obtain the solution of the rod problem on the triangular HL [as defined in Fig. 1(b)] following the same lines as in the square HL case discussed in Sec. II. The main differences are that now we have three lattice directions — which will be labeled as $A$, $B$ and $C$ here — and two sub-trees are attached to each of the two non-rooted sites of a rooted triangle, in order to obtain a lattice with coordination $q=6$; see Figs. 1(b) and 9. Since there are three lattice directions, but only two edges arriving at a root site, at first, one should distinguish among the situations where such edges are of type $AB$, $BC$ or $CA$. To keep the notation as simple as possible, we will name the ppf’s according to the label of the direction opposite (i.e., not connecting) to the root site. Thereby, for the configurations with no reaching rod (from above) in the root site, we will have the ppf’s $a_{0}$, $b_{0}$ and $c_{0}$. In the same way, we shall have the ppf’s $b_{(i,A)}$ and $c_{(i,A)}$; $a_{(i,B)}$ and $c_{(i,B)}$; and $a_{(i,C)}$ and $b_{(i,C)}$, when the reaching rod has $i$ monomers incorporated to it in direction $A$, $B$ or $C$. By defining the sums:
@@
Considering the simplified solution, the partition function of the system, obtained by attaching six sub-trees to the three vertices of the central triangle, reads:
@@
and, then, the total density is $\rho=\rho_{A}+\rho_{B}+\rho_{C}$. Following Ref. drs11 , the order parameter for the isotropic-nematic transition will be defined as
@@
As demonstrated, e.g., in Ref. furlanALG , the bulk free energy per site, $\phi_{b}$, for the triangular HL is

### v2

@@ Appendix A Solution of the model on the Husimi lattice built with triangles
We may obtain the solution of the rod problem on the triangular HL [as defined in Fig. 1(b)] following the same lines as in the square HL case discussed in Sec. 2. The main differences are that now we have three lattice directions — which will be labeled as $A$, $B$ and $C$ here — and two sub-cactuses are attached to each of the two non-rooted sites of a rooted triangle, in order to obtain a lattice with coordination $q=6$; see Figs. 1(b) and 9. Since there are three lattice directions, but only two edges arriving at a root site, at first, one should distinguish among the situations where such edges are of type $AB$, $BC$ or $CA$. To keep the notation as simple as possible, we will name the ppf’s according to the label of the direction opposite (i.e., not connecting) to the root site. Thereby, for the configurations with no reaching rod (from above) in the root site, we will have the ppf’s $a_{0}$, $b_{0}$ and $c_{0}$. In the same way, we shall have the ppf’s $b_{(i,A)}$ and $c_{(i,A)}$; $a_{(i,B)}$ and $c_{(i,B)}$; and $a_{(i,C)}$ and $b_{(i,C)}$, when the reaching rod has $i$ monomers incorporated to it in direction $A$, $B$ or $C$. By defining the sums:
@@
Considering the simplified solution, the partition function of the system, obtained by attaching six sub-cactuses to the three vertices of the central triangle, reads:
@@
and, then, the total density is $\rho=\rho_{A}+\rho_{B}+\rho_{C}$. Following Ref. [37], the order parameter for the isotropic-nematic transition will be defined as
@@
As demonstrated, e.g., in Ref. [45], the bulk free energy per site, $\phi_{b}$, for the triangular HL is